OpenStudy (anonymous):

What is the value of z so that -9 and 9 are both solutions of x^2 + z = 103 ? A ) -22 B ) 3 C ) 22 D ) 184 I think the answer is A?

OpenStudy (anonymous):

None of these are correct.. No matter what, you're eventually going to end up with a negative and a positive in the equation, and that can't sum up to a positive..

OpenStudy (anonymous):

Can you take a picture of the original sheet and show it to me?

OpenStudy (anonymous):

Is this from Icademy?

OpenStudy (anonymous):

Yes, I believe the answer is either 22 or -22 but I'm not sure which?

OpenStudy (anonymous):

Wait I think it's C ) 22! :)

OpenStudy (anonymous):

-81+22= -59. That can't be it.

OpenStudy (anonymous):

x^2 + 22 = 103 +-22 x^2 = 81 Then you get the square root of 81 which is 9 or -9?

OpenStudy (anonymous):

The square root of -9 is -81 though, so it wouldn't be a positive answer like 103. This question is just weird.

OpenStudy (anonymous):

In my grade we aren't dealing with negative square roots. Also, -9 x -9 would be 81? Haha I agree with you on that! I just did the Quick Check and it was 22, though.

OpenStudy (anonymous):

For the positive nine, Z would be 22. For the negative nine, Z would be 184

OpenStudy (anonymous):

Okay, I was doing it in my head for the square root. Sorry haha.

OpenStudy (anonymous):

It's fine. Thank you for the help :)

OpenStudy (anonymous):

x^2=103-z x=sqrt(103-z) 9=sqrt(103-z) squaring both side 81=103-z z=103-81=22